# LLM Personas as a Substitute for Field Experiments in Method Benchmarking

### v1

@@ 4.1 Literal panel change iff just panel change (JPC)
Consider two benchmarks $B$ and $B^{\prime}$ on the same artifact space $\mathcal{W}$ and feedback space $\mathcal{O}$. We say swapping $B$ for $B^{\prime}$ is just panel change (JPC) on the method interface if there exist Markov kernels $Q:\mathcal{W}\mapsto\Delta\mathcal{O}$ and $Q^{\prime}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that, for every algorithm $\mathcal{A}$ and every horizon $T$, the observable transcript laws factorize as
@@
Definition 2 is intentionally method-centric: it quantifies over all (possibly adaptive) algorithms and asks whether, from their perspective, the benchmark is an oracle channel that depends only on the current submission $w_{t}$, with no additional side channels or algorithm-dependent behavior. Crucially, JPC does not require $Q=Q^{\prime}$ (scores need not match across humans and personas); it only requires that the form of the interaction is preserved and that the swap can be summarized entirely by replacing one artifact-to-score kernel by another. This is precisely the sense in which swapping evaluators should resemble a “panel change” rather than a change in the rules of the game; the following Lemma 4.1 and 4.2 formalize the mathematical equivalence.
@@
Assume Aggregate-only observation (AO), i.e., the algorithm observes only $o_{t}\in\mathcal{O}$ each round. Let $B=(P,I,\Gamma,L)$ and $B^{\prime}=(P^{\prime},I,\Gamma,L)$ differ by a literal panel change in the sense of Definition 1. Define the reduced-form kernels
@@
Suppose two benchmarks $B$ and $B^{\prime}$ satisfy JPC in the sense of Definition 2, i.e., there exist Markov kernels $Q,Q^{\prime}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every algorithm $\mathcal{A}$ and horizon $T$,
@@
Then there exists a pair of benchmarks $\overline{B}=(\overline{P},\overline{I},\overline{\Gamma},\overline{L})$ and $\overline{B}^{\prime}=(\overline{P}^{\prime},\overline{I},\overline{\Gamma},\overline{L})$ that differ by a literal panel change (i.e., only $\overline{P}$ is replaced by $\overline{P}^{\prime}$), such that for every algorithm $\mathcal{A}$ and horizon $T$,
@@
Lemma 4.1 says that, if you literally change only the panel distribution $P$ in the panel–instrument–aggregation model, then the observable transcript laws change only by replacing the reduced-form kernel $Q_{P,I}$ by $Q_{P^{\prime},I}$, i.e., JPC holds. Lemma 4.2 says the converse at the interface level: any JPC swap can be realized exactly as a literal panel change in a (possibly abstract) benchmark representation. Thus Definition 2 is mathematically equivalent to “literal panel change” as an interface-level notion (i.e., up to equality of transcript laws for every adaptive algorithm).
@@
4.2 Just panel change (JPC) iff (AO)+(AB)
Lemma 4.3 ((JPC) $\iff$ (AO)+(AB)).
@@ 4.2 Just panel change (JPC) iff (AO)+(AB)
    (AO)+(AB). The protocol satisfies aggregate-only observation (AO) (Assumption 1) and each benchmark is algorithm-blind (AB) (Assumption 2).
At a technical level, JPC is defined by a factorization of the transcript laws through an artifact-only kernel. This makes the main theorem appear almost tautological: if we assume conditions that guarantee exactly such a factorization, we recover JPC. The point of stating the equivalence is not to rebrand a definition, but to (i) express JPC in terms of two concrete and auditable benchmark-hygiene requirements (what the benchmark reveals, and whether it is provenance-blind), and (ii) cleanly separate the two ways persona benchmarking can fail as a substitute for field experiments: leakage of micro-level information (violating AO) and algorithm/provenance dependence (violating AB).
@@
If (AB) fails, the benchmark is not a well-defined oracle environment.
When (AB) fails, there is no single kernel $Q(\cdot\mid w)$ that governs the returned score across algorithms. Equivalently, the evaluator’s behavior depends on provenance/identity or on interaction history in a way that is not summarized by the submitted artifact. In such a case, the benchmark cannot be treated as an artifact-only oracle channel, so “panel change” is not an identified description of the swap. See Appendix B.2 for when AB can be justified.
@@ 5 Beyond validity: When does a persona panel constitute an useful benchmark?
As we discussed in section 4, (AO)+(AB) characterizes when and only when persona benchmarking is a valid benchmark, satisfying (JPC). But validity, which is an identification argument, is not necessarily equivalent to usefulness. In this section, we formalize when and when not a persona-based LLM simulation can be a perfectly valid, but much less useful, benchmark compared to field experiments. We show that 1) dataset size is what really matters for a persona-based benchmark, or any benchmark, to be considered a useful benchmark, and 2) how to find such a required dataset size empirically.
As a starting point, recall that, under (AO)+(AB), the method interacts with the persona benchmark only through the reduced-form kernel
@@ 5.2 Formalization of usefulness: discriminability
Lemma 5.2 states that $\kappa_{Q}$ can be understood as the “per-data” information available to distinguish two meaningfully different artifacts (separated by at least $r$). When $\kappa_{Q}$ is small, because the benchmark is flat (small $\Delta$) or noisy (large $\sigma^{2}$), then any algorithm requires many independent data samples to reliably decide whether one artifact is better than another. That is, this scaling clarifies the effective size of dataset for benchmarking.
@@ 5.4 Choice of $r$ and $d_{\mathcal{W}}$
This makes $r$ interpretable as a step size in the space the algorithm explores. Examples include the scaled Euclidean distance on continuous hyperparameters or the edit distance on a structured prompt template.
@@
  • Model or policy variants: set $r$ to the smallest recipe change you would treat as a distinct method (e.g., one additional fine-tuning epoch, one dataset mixture adjustment above a threshold, a decoding rule change). If using a behavioral distance, $r$ can be a small divergence level that corresponds to a “barely different” behavior on the reference model (calibrated by pilot runs).
@@ 6 Conclusion
We characterized when LLM-persona panels can substitute for human field experiments as a benchmark interface for method development. Our main result shows that persona vs. human evaluation is just panel change from the method’s perspective if and only if two benchmark-hygiene conditions hold: (i) aggregate-only observation (AO) and (ii) algorithm-blind evaluation (AB). When either fails, the benchmark can leak exploitable information or depend on provenance, breaking the interface-level equivalence.
We also separated validity from usefulness. A valid persona benchmark may still be less informative than field experimentation if the induced aggregate channel is too flat or noisy. Our discriminability $\kappa_{Q}$ (worst-case KL separation at resolution $r$) yields the corresponding budget scaling for reliable comparisons, on the order of $\kappa_{Q}^{-1}\log(1/\delta)$. Thus, beyond enforcing (AO)+(AB), the practical requirement is a sufficient persona dataset.
@@ Appendix A Deferred theoretical discussions
Fix any algorithm $\mathcal{A}$ with submission kernels $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$ under (AO), and fix a round $t$. Under benchmark $B=(P,I,\Gamma,L)$, conditional on the submitted artifact $w_{t}$, the benchmark generates
@@
This is exactly the (AB)-type conditional independence statement with kernel $Q_{P,I}$. Combining this with the fact that $w_{t}\sim\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$ under (AO), the standard sequential composition / chain rule for Markov kernels yields, for every horizon $T$,
@@
Moreover, under $\overline{B}$, each benchmark call on $w$ samples $p=0$ a.s., then outputs $o\sim\overline{I}(\cdot\mid w,0)=Q(\cdot\mid w)$. Thus the reduced-form kernel of $\overline{B}$ is exactly $Q$. Similarly, the reduced-form kernel of $\overline{B}^{\prime}$ is exactly $Q^{\prime}$. Therefore, for every algorithm $\mathcal{A}$ and horizon $T$, the transcript laws under $\overline{B}$ and $\overline{B}^{\prime}$ satisfy the same factorizations as in the displayed JPC equations, which implies
@@
$(2)\Rightarrow(1)$: (AO)+(AB) imply JPC. Assume (AO) holds and each benchmark is algorithm-blind (AB) in the sense of Assumptions 12. Apply Lemma A.1 to the human benchmark $B_{\mathrm{hum}}$ to obtain a kernel $Q_{\mathrm{hum}}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every algorithm $\mathcal{A}$ and horizon $T$,
@@
$(1)\Rightarrow(2)$: JPC implies (AO)+(AB). Assume the swap is JPC in the sense of Definition 2. By that definition, the transcript laws of both benchmarks factorize for every algorithm $\mathcal{A}$ and every horizon $T$ through some kernels $Q$ and $Q^{\prime}$.
First, (AO) is the observation/information structure assumed in Definition 2 (the algorithm’s interaction is summarized by the aggregate-history $\mathcal{I}_{t-1}$ and transcript $(w_{t},o_{t})$). Second, applying Lemma A.2 to $B_{\mathrm{hum}}$ and $B_{\mathrm{pers}}$ shows that each benchmark satisfies (AB) with kernels $Q_{\mathrm{hum}}=Q$ and $Q_{\mathrm{pers}}=Q^{\prime}$ respectively.
Therefore, (AO)+(AB) hold. ∎
Lemma A.1 (Transcript factorization under (AO)+(AB)).
Fix a benchmark $B$ with artifact space $\mathcal{W}$ and feedback space $\mathcal{O}$. Assume (AO) and (AB) hold for $B$, i.e., there exists a Markov kernel $Q_{B}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every algorithm $\mathcal{A}$, every round $t$, and every measurable $A\subseteq\mathcal{O}$,
@@
Then for every algorithm $\mathcal{A}$ (with submission kernels $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$ under (AO)) and every horizon $T$,
@@
Fix an algorithm $\mathcal{A}$ and a round $t$.
@@
Step 2. Under (AO), the algorithm selects $w_{t}$ using the kernel $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$. Combining with Step 1, conditional on $\mathcal{I}_{t-1}$ we can generate $(w_{t},o_{t})$ by:
@@
Lemma A.2 (Factorization implies (AB)).
Fix a benchmark $B$. Suppose that under (AO), there exists a Markov kernel $Q_{B}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every algorithm $\mathcal{A}$ and every horizon $T$,
@@
Then $B$ satisfies (AB) in the sense of Assumption 2; namely, for every algorithm $\mathcal{A}$, every $t$, and measurable $A\subseteq\mathcal{O}$,
@@
since $Q_{B}(A\mid w_{t})$ is $\sigma(w_{t})$-measurable. This is exactly (AB). ∎
@@ B.1 AO and JPC - a counterexample
The equivalence in Lemma 4.3 highlights that (AO) is not merely technical: if micro-level information leaks (raw votes, rater identities, ordering, etc.), then two evaluation pipelines can induce the same aggregate kernel $Q(\cdot\mid w)$ yet still be distinguishable (and exploitable) by an adaptive algorithm. The following minimal construction makes this necessity direction concrete.
@@
There exist two benchmarks $B$ and $B^{\prime}$ that induce the same reduced-form kernel $Q(\cdot\mid w)$ on the aggregate feedback space $\mathcal{O}$, but such that if the benchmark reveals the raw vote vector $(Z_{1},\dots,Z_{L})$ to the algorithm (violating (AO)), then there is an adaptive algorithm $\mathcal{A}$ whose aggregate transcript law $\mathbb{P}^{\mathcal{A}}(w_{1:T},o_{1:T})$ differs between $B$ and $B^{\prime}$. Consequently, the swap cannot be “just panel change” on the method interface as in Definition 2.
@@
We specify two benchmarks and an adaptive algorithm.
@@
Now suppose (AO) is violated and the benchmark releases the raw vote vector $(Z_{1},Z_{2})$ (or $(Z^{\prime}_{1},Z^{\prime}_{2})$) to the algorithm in addition to the aggregate $o$. Consider the following horizon-$T=2$ adaptive algorithm $\mathcal{A}$:

### v2

@@ 4.1 Literal panel change iff just panel change (JPC)
Consider two benchmarks $B$ and $B^{\prime}$ on the same artifact space $\mathcal{W}$ and feedback space $\mathcal{O}$. We say swapping $B$ for $B^{\prime}$ is just panel change (JPC) on the method interface if there exist Markov kernels $Q:\mathcal{W}\mapsto\Delta\mathcal{O}$ and $Q^{\prime}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that, for every method $\mathcal{A}$ and every horizon $T$, the observable transcript laws factorize as
@@
Definition 2 is intentionally method-centric: it quantifies over all (possibly adaptive) methods and asks whether, from their perspective, the benchmark is an oracle channel that depends only on the current submission $w_{t}$, with no additional side channels or method-dependent behavior. Crucially, JPC does not require $Q=Q^{\prime}$ (scores need not match across humans and personas); it only requires that the form of the interaction is preserved and that the swap can be summarized entirely by replacing one artifact-to-score kernel by another. This is precisely the sense in which swapping evaluators should resemble a “panel change” rather than a change in the rules of the game; the following Lemma 4.1 and 4.2 formalize the mathematical equivalence.
@@
Assume Aggregate-only observation (AO), i.e., the method observes only $o_{t}\in\mathcal{O}$ each round. Let $B=(P,I,\Gamma,L)$ and $B^{\prime}=(P^{\prime},I,\Gamma,L)$ differ by a literal panel change in the sense of Definition 1. Define the reduced-form kernels
@@
Suppose two benchmarks $B$ and $B^{\prime}$ satisfy JPC in the sense of Definition 2, i.e., there exist Markov kernels $Q,Q^{\prime}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every method $\mathcal{A}$ and horizon $T$,
@@
Then there exists a pair of benchmarks $\overline{B}=(\overline{P},\overline{I},\overline{\Gamma},\overline{L})$ and $\overline{B}^{\prime}=(\overline{P}^{\prime},\overline{I},\overline{\Gamma},\overline{L})$ that differ by a literal panel change (i.e., only $\overline{P}$ is replaced by $\overline{P}^{\prime}$), such that for every method $\mathcal{A}$ and horizon $T$,
@@
Lemma 4.1 says that, if you literally change only the panel distribution $P$ in the panel–instrument–aggregation model, then the observable transcript laws change only by replacing the reduced-form kernel $Q_{P,I}$ by $Q_{P^{\prime},I}$, i.e., JPC holds. Lemma 4.2 says the converse at the interface level: any JPC swap can be realized exactly as a literal panel change in a (possibly abstract) benchmark representation. Thus Definition 2 is mathematically equivalent to “literal panel change” as an interface-level notion (i.e., up to equality of transcript laws for every adaptive method).
@@
4.2 Just panel change (JPC) iff (AO)+(MB)
Lemma 4.3 ((JPC) $\iff$ (AO)+(MB)).
@@ 4.2 Just panel change (JPC) iff (AO)+(MB)
    (AO)+(MB). The protocol satisfies aggregate-only observation (AO) (Assumption 1) and each benchmark is method-blind (MB) (Assumption 2).
At a technical level, JPC is defined by a factorization of the transcript laws through an artifact-only kernel. This makes the main theorem appear almost tautological: if we assume conditions that guarantee exactly such a factorization, we recover JPC. The point of stating the equivalence is not to rebrand a definition, but to (i) express JPC in terms of two concrete and auditable benchmark-hygiene requirements (what the benchmark reveals, and whether it is provenance-blind), and (ii) cleanly separate the two ways persona benchmarking can fail as a substitute for field experiments: leakage of micro-level information (violating AO) and method/provenance dependence (violating MB).
@@
If (MB) fails, the benchmark is not a well-defined oracle environment.
When (MB) fails, there is no single kernel $Q(\cdot\mid w)$ that governs the returned score across methods. Equivalently, the evaluator’s behavior depends on provenance/identity or on interaction history in a way that is not summarized by the submitted artifact. In such a case, the benchmark cannot be treated as an artifact-only oracle channel, so “panel change” is not an identified description of the swap. See Appendix B.2 for when MB can be justified.
@@ 5 Beyond validity: When does a persona panel constitute an useful benchmark?
As we discussed in section 4, (AO)+(MB) characterizes when and only when persona benchmarking is a valid benchmark, satisfying (JPC). But validity, which is an identification argument, is not necessarily equivalent to usefulness. In this section, we formalize when and when not a persona-based LLM simulation can be a perfectly valid, but much less useful, benchmark compared to field experiments. We show that 1) dataset size is what really matters for a persona-based benchmark, or any benchmark, to be considered a useful benchmark, and 2) how to find such a required dataset size empirically.
As a starting point, recall that, under (AO)+(MB), the method interacts with the persona benchmark only through the reduced-form kernel
@@ 5.2 Formalization of usefulness: discriminability
Lemma 5.2 states that $\kappa_{Q}$ can be understood as the “per-data” information available to distinguish two meaningfully different artifacts (separated by at least $r$). When $\kappa_{Q}$ is small, because the benchmark is flat (small $\Delta$) or noisy (large $\sigma^{2}$), then any method requires many independent data samples to reliably decide whether one artifact is better than another. That is, this scaling clarifies the effective size of dataset for benchmarking.
@@ 5.4 Choice of $r$ and $d_{\mathcal{W}}$
This makes $r$ interpretable as a step size in the space the method explores. Examples include the scaled Euclidean distance on continuous hyperparameters or the edit distance on a structured prompt template.
@@
  • Model or policy variants: set $r$ to the smallest recipe change you would treat as a distinct method (e.g., one additional fine-tuning epoch, one dataset mixture adjustment above a threshold, a decoding rule change).
@@
5.5 Experiments
As a proof-of-concept, Appendix C demonstrates how our discriminability-based analysis can be used to calibrate the persona dataset size needed for reliable method comparison. Concretely, we apply the procedure to the persona-simulation ad benchmark used to evaluate TextBO, a prompt-optimization based self-improving AI system (Kang and Yoganarasimhan, 2025), and compute the implied number of independent persona evaluations required to distinguish one-step prompt improvements at a chosen confidence level.
@@ 6 Conclusion
We characterized when LLM-persona panels can substitute for human field experiments as a benchmark interface for method development. Our main result shows that persona vs. human evaluation is just panel change from the method’s perspective if and only if two benchmark-hygiene conditions hold: (i) aggregate-only observation (AO) and (ii) method-blind evaluation (MB). When either fails, the benchmark can leak exploitable information or depend on provenance, breaking the interface-level equivalence.
We also separated validity from usefulness. A valid persona benchmark may still be less informative than field experimentation if the induced aggregate channel is too flat or noisy. Our discriminability $\kappa_{Q}$ (worst-case KL separation at resolution $r$) yields the corresponding budget scaling for reliable comparisons, on the order of $\kappa_{Q}^{-1}\log(1/\delta)$. Thus, beyond enforcing (AO)+(MB), the practical requirement is a sufficient persona dataset.
@@ Appendix A Deferred theoretical discussions
Fix any method $\mathcal{A}$ with submission kernels $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$ under (AO), and fix a round $t$. Under benchmark $B=(P,I,\Gamma,L)$, conditional on the submitted artifact $w_{t}$, the benchmark generates
@@
This is exactly the (MB)-type conditional independence statement with kernel $Q_{P,I}$. Combining this with the fact that $w_{t}\sim\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$ under (AO), the standard sequential composition / chain rule for Markov kernels yields, for every horizon $T$,
@@
Moreover, under $\overline{B}$, each benchmark call on $w$ samples $p=0$ a.s., then outputs $o\sim\overline{I}(\cdot\mid w,0)=Q(\cdot\mid w)$. Thus the reduced-form kernel of $\overline{B}$ is exactly $Q$. Similarly, the reduced-form kernel of $\overline{B}^{\prime}$ is exactly $Q^{\prime}$. Therefore, for every method $\mathcal{A}$ and horizon $T$, the transcript laws under $\overline{B}$ and $\overline{B}^{\prime}$ satisfy the same factorizations as in the displayed JPC equations, which implies
@@
$(2)\Rightarrow(1)$: (AO)+(MB) imply JPC. Assume (AO) holds and each benchmark is method-blind (MB) in the sense of Assumptions 12. Apply Lemma A.1 to the human benchmark $B_{\mathrm{hum}}$ to obtain a kernel $Q_{\mathrm{hum}}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every method $\mathcal{A}$ and horizon $T$,
@@
$(1)\Rightarrow(2)$: JPC implies (AO)+(MB). Assume the swap is JPC in the sense of Definition 2. By that definition, the transcript laws of both benchmarks factorize for every method $\mathcal{A}$ and every horizon $T$ through some kernels $Q$ and $Q^{\prime}$.
First, (AO) is the observation/information structure assumed in Definition 2 (the method’s interaction is summarized by the aggregate-history $\mathcal{I}_{t-1}$ and transcript $(w_{t},o_{t})$). Second, applying Lemma A.2 to $B_{\mathrm{hum}}$ and $B_{\mathrm{pers}}$ shows that each benchmark satisfies (MB) with kernels $Q_{\mathrm{hum}}=Q$ and $Q_{\mathrm{pers}}=Q^{\prime}$ respectively.
Therefore, (AO)+(MB) hold. ∎
Lemma A.1 (Transcript factorization under (AO)+(MB)).
Fix a benchmark $B$ with artifact space $\mathcal{W}$ and feedback space $\mathcal{O}$. Assume (AO) and (MB) hold for $B$, i.e., there exists a Markov kernel $Q_{B}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every method $\mathcal{A}$, every round $t$, and every measurable $A\subseteq\mathcal{O}$,
@@
Then for every method $\mathcal{A}$ (with submission kernels $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$ under (AO)) and every horizon $T$,
@@
Fix an method $\mathcal{A}$ and a round $t$.
@@
Step 2. Under (AO), the method selects $w_{t}$ using the kernel $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$. Combining with Step 1, conditional on $\mathcal{I}_{t-1}$ we can generate $(w_{t},o_{t})$ by:
@@
Lemma A.2 (Factorization implies (MB)).
Fix a benchmark $B$. Suppose that under (AO), there exists a Markov kernel $Q_{B}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every method $\mathcal{A}$ and every horizon $T$,
@@
Then $B$ satisfies (MB) in the sense of Assumption 2; namely, for every method $\mathcal{A}$, every $t$, and measurable $A\subseteq\mathcal{O}$,
@@
since $Q_{B}(A\mid w_{t})$ is $\sigma(w_{t})$-measurable. This is exactly (MB). ∎
@@ B.1 AO and JPC - a counterexample
The equivalence in Lemma 4.3 highlights that (AO) is not merely technical: if micro-level information leaks (raw votes, rater identities, ordering, etc.), then two evaluation pipelines can induce the same aggregate kernel $Q(\cdot\mid w)$ yet still be distinguishable (and exploitable) by an adaptive method. The following minimal construction makes this necessity direction concrete.
@@
There exist two benchmarks $B$ and $B^{\prime}$ that induce the same reduced-form kernel $Q(\cdot\mid w)$ on the aggregate feedback space $\mathcal{O}$, but such that if the benchmark reveals the raw vote vector $(Z_{1},\dots,Z_{L})$ to the method (violating (AO)), then there is an adaptive method $\mathcal{A}$ whose aggregate transcript law $\mathbb{P}^{\mathcal{A}}(w_{1:T},o_{1:T})$ differs between $B$ and $B^{\prime}$. Consequently, the swap cannot be “just panel change” on the method interface as in Definition 2.
@@
We specify two benchmarks and an adaptive method.
@@
Now suppose (AO) is violated and the benchmark releases the raw vote vector $(Z_{1},Z_{2})$ (or $(Z^{\prime}_{1},Z^{\prime}_{2})$) to the method in addition to the aggregate $o$. Consider the following horizon-$T=2$ adaptive method $\mathcal{A}$: